WinArrange

Full documentation and hints for WinArrange under AmigaOS.

Window sorting tool for the active screen. 10 April 1995

Thomas Alexnat

Versions

Release of WinArrange:

1.0

WinArrange:

0.67 (02.03.95)

english documentation:

0.1 (21.02.95)

Operating system:

at least AmigaOS 2.04 (AmigaOS 3.1 is highly recommended!)

Copyright

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1995 Thomas Alexnat

To contact the author of all files which are involved with WinArrange (i.e. executable of WinArrange itself, the different documentations and some other relevant files), try to reach one of the following addresses:

         Thomas Alexnat
         Westendstrasse 7
         82110 Germering (Germany)
         Fido:     2:2480/700.5
         MAX:      91:5150/10.5
         E-Mail:   alexnat@eikon.e-technik.tu-muenchen.de

Please read on for more information about using WinArrange.

1 Purpose

WinArrange is a small utility to rearrange your windows on the active screen.

1.1 Why WinArrange ?

In former days I went crazy, while sorting and arranging windows on the workbench screen. Serveral shell, editor and normal workbench windows were untidying my beautiful workenvironment. I wasn’t really cabable of knowing what was really going on. The only solution for this problem was to rearrange all windows on this screen. Unfortunally, this event took me a lot of time. After rebooting and starting my programms my workbench was again in a mess. As you know time is money, I started to write this little programm to make the life under AmigaOS with Intuition windows and screens peacefuller. Another important point is, that I wanted to implement a feature, which nearly every operating system should support: The rearrangement of windows for accessing faster the applications.

1.2 What does it do ?

WinArranges has got the abillity to reorganzie the windows on the active screen. This can be done in many different ways. You can cascade, shuffle or sort the windows e.g. depending relatively on the current screen resolution or aspect ratio. If you don’t want to have the windows sorted, which are owned by the Workbench’s task, you can just switch this off. Here is a list of some features:

• many different sorting methods: e.g. horizontal, shuffle, chess, user-defined

• abillity to ignore windows for sorting (e.g. workbench window)

• executable since AmigaOS 2.04 (37.175)

• nearly all features AmigaOS provides are supported

• continuing support and development

2 Installation

At first you must check out, which version of AmigaOS you have installed. Ususally it should be something like 2.1 oder higher. Otherwise update your Hardware, to enjoy this litte utility. Installing WinArrange is very easy. There are two possibillities of making WinArrange start-ready for your system:

1. Double-click on the Install_English icon (or the desired language). For novice Users, the installation-script auto-configures WinArrange. For detailed information read the concerning chapters. The installation-procedure will also copy files like the onlinedocumentation and texinfofiles to the desired directories.

2. Expert-Users can also drop the WinArrange icon into the desired drawer of your workstation, due to the fact, that WinArrange is not too complicated.

3 Usage

This chapter will introduce you, dealing with WinArrange. See these few pages as a tutorial. You will learn how to start and configure WinArrange for your special own needs and wishes. Please make sure, that you have followed the installing instruction in the corresponding chapter and install script.

3.1 Tutorial

If you want a fast image, what WinArrange can do for you, you may double click on the WinArrange icon. The default setting is, that all resizable Windows on the active screen are sorted vertically.

3.1.1 Choosetype

If you don’t want, that the WorkBench’s windows (which are containing all the icons) are sorted, you can switch this off by adding/changing the tooltype CHOOSETYPE. Changing this tooltype is very easy:

1. Open the drawer on the WorkBench, where WinArrange is located.

2. Select the WinArrange Icon (just click once, not twice)

3. Hold down the menu button on your mouse (usually the right mouse button), and select in

    the menu ICON the item INFORMATION....

4. Now you can see, if already entered, different tooltypes.

5. Click on the line with the entry (CHOOSETYPE=NOWB).

6. Then remove the two brackets, or if the line doesn’t exist, add the tooltype CHOOSETYPE=NOWB also without the brackets and save these settings.

Now try to doubleclick on the icon. You will recognize, that WinArrange won’t sort the Work-Bench’s windows.

3.1.2 Sorttype

The default sorting method of WinArrange is very boring. You can try to use another system: CHESS — You have only to change to tooltype SORTTYPE from (SORTTYPE=HORIZ) to SORTTYPE=CHESS. Please proceed in the same way a showed above.

Now you have seen, what WinArrange can do for you. We’ll start now, with the detailed documentation

3.2 Shell-Options

As you have noticed, WinArrange is able to be started via a command shell or the AmigaOS WorkBench. At first I explain you the CLI-template-string. Open a command shell and go to the directory where WinArrange is located (e.g. double-click the Shell Icon in SYS:System/ and type

in CD Utilities):

To get the template-string^{1} type the following sequence:

     WinArrange ? [RETURN]

This should give you the follwing result:

     H=Horiz/K/N,V=Vert/K/N,C=Chess/S,S=Shuffle/S,WB=WBOnly/S,NoWB/S,Min/S,Max/S:

H=Horiz/K/N The Horiz parameter requires an integer number. In other words, it is are percentage number, which means, that in what way the windows are expanded or shrinked horiontally.

V=Vert/K/N

Same parameters and effects like Horiz, but the resizing affects the vertical size.

C=Chess/S This is a mutal exclude switch. All windows are sorted like a chess table. This method is very interesting for big resulutions and many windows. You can keep track of all your windows while seeing a part of the contents in the windows. The Horiz and the

1 The template-string should help the user a little bit. It displays all arguments and options, which are needed to start succsessfully a program. Usually this string can be accessed by adding a question mark `?' to the command.

Vert have got the highest priority, so if such a parameter is set, the Chess sorting method will be overiden by a normal Shuffle system.

S=Shuffle/S This is a mutal exclude switch. The windows are here placed in a diagonal line, beginning at the left side of the current screen. The top side and the window step depends on the active screenfont.

WB=WBOnly/S Using this switch will affect all windows which are owned by the workbench task. WinArrange checks if the workbench flag is set. Other windows are ignored. The opposite flag is NoWB

NoWB/S All windows which are not owned by the wokbench are involved in rearranging. The opposite command is WBOnly.

Min/S The Min switch is some kind of a macro. It sets the Horiz and Vert options to 1 percent. The Min command can be seen as a global iconify option. The opposite function is Max

Max/S This option sets Horiz and Vert to the maximum. The result is the opposite side of Min, as described above.

3.3 Configuring

As you have seen, WinArrange is a powerfull tool. To keep track of all windows and sorting methods, WinArrnge has got the abillity to be configured via a config-file. This file is stored in the logically drive ENV: and is named WinArrange.config.

3.3.1 Structure of the config-file

The config-file uses normal ASCII strings and tags and is also readable humans. Here follow some declarations which are needed to use the config-file.

Semicolons Semicolons ’;’ are treated like in the AmigaOS shell or Basic commands as remarks. A remark should start always at the beginning of a new line. Multiple remarks in a line are not allowed.

other conditions Please note, that WinArrange hasn’t got a professional config-file parser. If something weired happens, feel free to contact me.

To demomstrate, how the config-file works, I’ve included an example configuration-file.

; $VER: WinArrange.config 0.1 (20.07.94)
;  1994 by Thomas Alexnat
;
; This is a configfile for WinArrange
;
; 1.
;   Please note, that the keyword 'configtype', is a identifier
;   for WinArranges' configparser. It may happen, that the format
;   of the config changes in future. At the moment, the is only existing
;   one type: 1
;
; 2.
;   Semicolons ';' are treated like in AmigaOS or Basic as remarks
;
; 3.
;   It is also allowed to enter are blank line
;
; 4.
;   At this point of time, you may not use linebreaks. This means, that a
;   name has got to be at maximum, only a full line, preceeded by a normal
;   AmigaOS linebreak command.
;
; 5.
;   The config file should be plain AmigaOS-ASCII
;
; Thank you, for reading all this information. You may also cut out these
; remarks for speed improvisation
;Global options:
WBONLY FALSE

;Now all options and patterns for the windows:
ACCEPTWINDOW ToolmanagerDock
REJECTWINDOW WorkBench
;Now all options and patterns for the tasks:
ACCEPTTASK WorkBench
REJECTTASK CEd

3.4 Hints

Due to the fact, that WinArrange is not a AmigaOS Commodity, it hasn’t got the abillity to be started or activated via hotkeys. I recommend you to assign some keys with e.g. ToolManager, DockTool.

WinArrange prints always his version string out. If you don’t like that, just copy the output to a file (e.g. NIL:). To keep in control with your system, you may also enter something like this: WinArrange >LOGDIR:WinArrange.log. If WinArrange has got some problems, you’ll see what WinArrange dumped in his logfile.

Concept Index 8

Concept Index

C

Choosetype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 config-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Configuring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

F

fontsensitiv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

H

Hints and Tips. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

I

Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

P

Purpose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

S

Shell-Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Sorttype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Structure of the config-file. . . . . . . . . . . . . . . . . . . . . . . . . . . 5

T

Tutorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

U

Usage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Table of Contents

1 Purpose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Why WinArrange ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 What does it do ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3.1 Tutorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.1 Choosetype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.1.2 Sorttype . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3.2 Shell-Options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.3 Configuring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.3.1 Structure of the config-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.4 Hints. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8